# Sequential Recommendation for Optimizing Both Immediate Feedback and Long-term Retention

### v1

@@ Abstract.
In the landscape of Recommender System (RS) applications, reinforcement learning (RL) has recently emerged as a powerful tool, primarily due to its proficiency in optimizing long-term rewards. Nevertheless, it suffers from instability in the learning process, stemming from the intricate interactions among bootstrapping, off-policy training, and function approximation. Moreover, in multi-reward recommendation scenarios, designing a proper reward setting that reconciles the inner dynamics of various tasks is quite intricate. In response to these challenges, we introduce DT4IER, an advanced decision transformer-based recommendation model that is engineered to not only elevate the effectiveness of recommendations but also to achieve a harmonious balance between immediate user engagement and long-term retention. The DT4IER applies an innovative multi-reward design that adeptly balances short and long-term rewards with user-specific attributes, which serve to enhance the contextual richness of the reward sequence ensuring a more informed and personalized recommendation process. To enhance its predictive capabilities, DT4IER incorporates a high-dimensional encoder, skillfully designed to identify and leverage the intricate interrelations across diverse tasks. Furthermore, we integrate a contrastive learning approach within the action embedding predictions, a strategy that significantly boosts the model’s overall performance. Experiments on three real-world datasets demonstrate the effectiveness of DT4IER against state-of-the-art Sequential Recommender Systems (SRSs) and Multi-Task Learning (MTL) models in terms of both prediction accuracy and effectiveness in specific tasks. The source code is accessible online to facilitate replication  [^1] https://anonymous.4open.science/r/DT4IER-5837.
@@
journalyear: 2024copyright: acmcopyrightbooktitle: Proceedings of the 47th ACM SIGIR conference on Recommender Systems (SIGIR ’24), July 14-18, 2024, Washington D.C., USA
@@ 1. Introduction
In today’s digital age, platforms spanning from social media to e-commerce have led to an explosion in the amount of information available online, underscoring the importance of efficient navigation tools (Aceto et al., 2020). Recommender Systems (RS) have emerged as a crucial technology in this realm, adeptly improving content suggestions and optimizing recommendations according to user interests inferred from their historical engagement. In recent years, researchers have proposed various methods for Recommender Systems, encompassing collaborative filtering  (Mooney and Roy, 2000), matrix factorization-based approaches  (Koren et al., 2009), and those powered by deep learning  (Cheng et al., 2016; Zhang et al., 2019). Among them, transformer-based models, notably BERT4Rec (Sun et al., 2019) and SASRec (Kang and McAuley, 2018) have risen to prominence, redefining the landscape of recommendation systems. It is believed that the strength of transformers lies in the attention mechanisms (Vaswani et al., 2017), which can dynamically capture the inherent dependencies in data, offering an accurate understanding of user patterns. This adaptability and precision make transformers well-suited for navigating the realm of user preferences inference and immediate feedback prediction.
In practice, a range of metrics such as clicks, likes, and ratings are widely used as indicators of user preferences. Though effective, these immediate metrics offer limited insights into the user’s lasting impression in the long term. In some cases, they can be misleading indicators of content quality (Wu et al., 2017; Yi et al., 2014). For instance, content with an eye-catching title but poor content quality may initially draw attention and obtain positive feedback, but only to later abuse the users’ trust and undermine their loyalty. This highlights a pressing need to balance the RS’s focus between short-term user engagement and long-term user satisfaction (Viljanen et al., 2016) . To this end, methods based on Reinforcement Learning (RL) have emerged (Chen et al., 2021b). By modeling sequential user behaviors using a Markov Decision Process (MDP), RL-based systems can dynamically adapt recommendations at every user interaction juncture (Zhao et al., 2011; Mahmood and Ricci, 2007). The strength of RL lies in its capability to optimize cumulative rewards over extended sequences in the future, allowing it to capture long-term objectives and trends, and ensure sustained user engagement. A recent finding (Cai et al., 2023a) shows that RL-based RS can even directly optimize user retention, a crucial metric often ignored but paramount in real-world business contexts (Zou et al., 2019) that is closely related to daily active users (DAU).
@@
  • Our innovative framework applies a novel multi-reward setting that balances immediate user responses with long-term retention signals by user-specific features, and then complements by a corresponding high-dimensional embedding module and a contrastive loss term.
@@ 3.2. Framework Overview
  • Embedding Module: The modified trajectory data then undergoes an embedding transformation to be represented as dense vectors. To be specific, the state and action sequences are processed using a Gated Recurrent Unit (GRU) (Cho et al., 2014), detailed further in Appendix D.1. In contrast, the reward sequence employs a high-dimensional encoder which is further explained in Section 3.4. Additionally, a positional session encoding is integrated.
  • Transformer Decision Block: Upon processing, the dense trajectory vectors serve as context, guiding the generation of action embeddings for the subsequent timestamp. This decision-making module is underpinned by the transformer model, detailed further in the Appendix D.2.
  • Action Decoding Block: Using the predicted action embeddings, our decoding unit strives to construct an action sequence that aligns closely with the reference of ground truth actions, detailed further in Appendix D.3.
@@ 3.3. Adaptive RTG Balancing
In settings that involve long short-term reward optimization, striking a balance between short-term and long-term performance metrics is notably challenging. This complexity arises from inherent conflicts between immediate and deferred objectives, as well as from the convoluted interdependencies among tasks (Ratner et al., 2018). To adeptly balance this equilibrium, we propose a novel strategy: the balancing of the RTG sequence, adaptively controlled by distinct user features. Since these user features hold a wealth of information about user preferences, they serve as a reliable guide for understanding user behaviors, bridging the gap between their immediate feedback and long-term patterns.
@@
    where $z_{i}$ represents the $i$-th element of the output vector $\mathbf{z}$, and $K$ is the total number of elements in $\mathbf{z}$, representing the different weights for balancing immediate response and long-term retention.
@@ 3.4. Multi-reward Embedding
    (11) $\widehat{\textbf{E}}^{R}=concate(w_{1}E^{R}_{1,t},w_{2}E^{R}_{2,t})$
@@ 4.1. Dataset
We carried out our experiment using three real-world recommendation datasets with details presented in Appendix B:
@@ 4.3. Baselines
  • RMTL (Liu et al., 2023): This is a reinforcement learning-based multi-task recommendation model with adaptive loss weights.
@@
Furthermore, we’ve adapted the model architectures of MMoE and PLE to facilitate sequential recommendations, utilizing a weighted score derived from multiple tasks. For more information, please refer to Appendix C.
@@ 4.5. Overall Performance and Comparison
  • Superiority of DT4IER: The DT4IER model consistently outperforms the four baselines in both the realms of recommendation accuracy and user retention score across datasets. Particularly on the Kuairand-Pure dataset, DT4IER exhibits an impressive 0.07-0.09 improvement in recommendation accuracy compared to the top-performing baseline. By interpreting RL as an autoregressive challenge and integrating a well-designed structure tailored for multiple rewards, our model strikes a balance between immediate feedback and long-term retention, resulting in significant enhancements in recommendation performance while retaining user engagement.
@@ 4.8. Case Study
In this subsection, we highlight DT4IER’s effectiveness in improving recommendation performance through a case study on the Kuairand-Pure dataset. The selected user for this study has a history of four interactions, with the ground truth action sequence comprising details of five different videos. Figure 2 shows that when provided with the user’s state and the goal of maximizing retention, DT4Rec recommends four videos. This recommendation achieves a BLEU score of 0.625 and an NDCG of 0.710. In contrast, DT4IER outperforms by recommending five actions, all of which align with the ground truth action sequence. This leads to a higher prediction accuracy, with a BLEU score of 0.887 and an NDCG of 0.947. These results underscore our model’s enhanced efficiency, attributed to its simultaneous consideration of short-term clicks and long-term retention information.
@@ 5.1. Sequential Recommender Systems
Sequential Recommendation refers to a recommendation system paradigm that models patterns of user behavior and items over time to suggest relevant products or content (Li et al., 2022). Among the myriad of approaches available, Markov Chains (MCs) (Rendle et al., 2010) and Recurrent Neural Networks (RNNs) stand out for their prowess in sequence modeling (Donkers et al., 2017). The adaptability and capability of RNNs in integrating diverse information make them especially effective for crafting sequential recommendations. RU4Rec (Tan et al., 2016) was a pioneer in leveraging RNNs, tapping into their potential to process sequential data. Subsequently, GRU4Rec (Hidasi et al., 2016) unveiled a parallel RNN structure to process item features, thus boosting recommendation quality. Further, the generalization power of the transformer architecture has given rise to its prominence in Sequential Recommendation Systems, birthing models such as BERT4Rec (Sun et al., 2019) and SASRec (Kang and McAuley, 2018). Building upon these foundations, the decision transformer was conceptualized to tackle user retention challenges, aiming to directly predict actions using a reward-driven autoregressive framework (Zhao et al., 2023b). However, a gap remains in current research concerning the optimization of multi-reward settings, a crucial aspect for a plethora of applications.
@@ 5.2. Reinforcement Learning Based Recommender Systems
The validity of applying RL-based solutions (Sutton and Barto, 2018; Afsar et al., 2021; Wang et al., 2022; Zhang et al., 2022) for sequential recommendation comes from the assumption of Markov Decision Process (Shani et al., 2005). And the key advantage of RL solutions is the ability to improve the expected cumulative reward of future interactions with users, rather than optimizing the one-step recommendation. Specifically, for scenarios with small recommendation spaces, one can use tabular-based (Mahmood and Ricci, 2007; Moling et al., 2012) or value-based methods  (Taghipour et al., 2007; Zheng et al., 2018; Zhao et al., 2018; Ie et al., 2019b) to directly evaluate the long-term value of the recommendation; For scenarios where action spaces are large, policy gradient methods (Sun and Zhang, 2018; Chen et al., 2019a, b) and actor-critic methods (Sutton et al., 1999; Peters and Schaal, 2008; Bhatnagar et al., 2007; Degris et al., 2012; Dulac-Arnold et al., 2015; Liu et al., 2018, 2020; Xin et al., 2020; Zhao et al., 2020; Cai et al., 2023b) are adopted to guide the policy towards better recommendation quality. Knowing that the web service may want to optimize multiple metrics, several works have discussed the challenge of multi-objective optimization (Chen et al., 2021c; Cai et al., 2023b) where the user behaviors might have different distributional patterns. Among all user feedback signals, user retention has been considered one of the most challenging to optimize, while recent work has shown a possible RL-based approach (Cai et al., 2023a) for this uphill struggle. This work aims to simultaneously optimize immediate feedback and user retention. Similar to our work, to overcome the gap between experiments on real user environments and offline evaluations, user simulators are widely used to bypass this paradox (Ie et al., 2019c; Zhao et al., 2019, 2023a). Our approach can be thought of as a likelihood-based method while employing a sequence modeling objective rather than relying on variational techniques.
@@ 5.3. Multi-task Learning in Recommender Systems
Multi-task learning (MTL) is a machine learning technique that addresses multiple tasks simultaneously  (Zhang and Yang, 2021). It captures a shared representation of the input through a shared bottom layer and then processes each task using distinct networks with task-specific weights. The overall performance is further enhanced by the knowledge transfer between tasks. This approach has become particularly popular in recommender systems, thanks to its prowess in effectively sharing data across tasks and in recognizing a range of user behaviors (Ma et al., 2018a; Lu et al., 2018; Hadash et al., 2018; Pan et al., 2019; Pei et al., 2019). A significant portion of recent research targets the enhancement of these architectures to promote more effective knowledge sharing. Notably, some innovations focus on introducing constraints to task-specific parameters (Duong et al., 2015; Misra et al., 2016; Yang and Hospedales, 2016), while others aim to clearly demarcate shared from task-specific parameters (Ma et al., 2018b; Tang et al., 2020). The principal goal of these strategies is to amplify knowledge transfer through improved feature representation. Additionally, some researchers are exploring the potential of reinforcement learning to enhance the MTL model by adjusting loss function weights (Liu et al., 2023). While many approaches prioritize item-wise modeling, our research delves into the challenges of balancing both short-term and long-term rewards in sequential recommendation settings. Furthermore, we have refined the methodologies of both the PLE and MMoE models to cater specifically to sequential recommendations.
@@ Appendix A HYPER-PARAMETER SELECTION
Appendix B DataSet Statistics
The statistics of three business datasets we use in this paper.
Table 3. Datasets Statistics
Dataset # Users # Items # Interactions # Density
Kuairand-Pure 27,285 7,551 1,436,609 0.70%
MovieLens-25M 162,000 62,000 25,000,000 0.25%
RetailRocket 1,407,580 234,843 2,684,433 0.0008%
Appendix C Inference setting for MMoE & PLE
Among MTL models, MMoE and PLE have consistently showcased outstanding performance, positioning them as standard choices for benchmarking against other approaches. In our research, we’ve fine-tuned their inference mechanisms to tailor them specifically for a sequential multi-task recommendation context. The adjustments we made are detailed below:
  • Similar to the Decision Transformer (DT) setup, we structured the dataset around user interactions within a specific day, treating each set of interactions as an individual session. Each of these sessions encompasses a maximum of 30 items, ensuring a manageable and consistent input size.
  • Our primary tasks of interest revolve around two crucial user behaviors: the click rate, which indicates immediate user engagement, and the user retention frequency, which serves as a proxy for long-term user loyalty and interest.
  • Once the 30 items are fed into the model, it predicts the click rate and retention frequency for each item. Following this, we employ a parameter $\eta$ to compute a weighted score for these two metrics. This score assists in ranking the items in descending order of relevance. Based on this refined list, we select the top 15 items. Subsequently, prediction metrics are computed by contrasting these selected items with the ground truth.
To ensure fairness in our evaluation, we carried out a grid search for the parameter $\eta$ within the range of 0.1 to 0.9. The optimal value emerged as 0.8, indicating a strong weighting towards the click rate. This observation sheds light on the notably high CTR precision observed for PLE and MMoE on the ML-25M dataset.
Appendix D Other Model details
This part details the state-action encoder, transformer decision block, and action decoder.
D.1. State-action Encoder

### v2

@@ Abstract.
In Recommender System (RS) applications, reinforcement learning (RL) has recently emerged as a powerful tool, primarily due to its proficiency in optimizing long-term rewards. Nevertheless, it suffers from instability in the learning process, stemming from the intricate interactions among bootstrapping, off-policy training, and function approximation. Moreover, in multi-reward recommendation scenarios, designing a proper reward setting that reconciles the inner dynamics of various tasks is quite intricate. To this end, we propose a novel decision transformer-based recommendation model, DT4IER, to not only elevate the effectiveness of recommendations but also to achieve a harmonious balance between immediate user engagement and long-term retention. The DT4IER applies an innovative multi-reward design that adeptly balances short and long-term rewards with user-specific attributes, which serve to enhance the contextual richness of the reward sequence, ensuring a more informed and personalized recommendation process. To enhance its predictive capabilities, DT4IER incorporates a high-dimensional encoder to identify and leverage the intricate interrelations across diverse tasks. Furthermore, we integrate a contrastive learning approach within the action embedding predictions, significantly boosting the model’s overall performance. Experiments on three real-world datasets demonstrate the effectiveness of DT4IER against state-of-the-art baselines in terms of both immediate user engagement and long-term retention. The source code is accessible online to facilitate replication  [^1] https://github.com/Applied-Machine-Learning-Lab/DT4IER.
@@
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3626772.3657829conference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAisbn: 979-8-4007-0431-4/24/07ccs: Information systems Recommender systems
@@ 1. Introduction
In today’s digital age, platforms spanning from social media to e-commerce have led to an explosion in the amount of information available online, underscoring the importance of efficient navigation tools (Aceto et al., 2020). Recommender Systems (RS) have emerged as a crucial technology in this realm, adeptly improving content suggestions and optimizing recommendations according to user interests inferred from their historical engagement. In recent years, researchers have proposed various methods for Recommender Systems, encompassing collaborative filtering  (Mooney and Roy, 2000), matrix factorization-based approaches  (Koren et al., 2009), and those powered by deep learning  (Cheng et al., 2016; Zhang et al., 2019; Chen et al., 2022). Among them, transformer-based models, notably BERT4Rec (Sun et al., 2019) and SASRec (Kang and McAuley, 2018) have risen to prominence, redefining the landscape of recommendation systems. It is believed that the strength of transformers lies in the attention mechanisms (Vaswani et al., 2017), which can dynamically capture the inherent dependencies in data, offering an accurate understanding of user patterns. This adaptability and precision make transformers well-suited for navigating the realm of user preferences inference and immediate feedback prediction.
In practice, a range of metrics such as clicks, likes, and ratings are widely used as indicators of user preferences. Though effective, these immediate metrics offer limited insights into the user’s lasting impression in the long term. In some cases, they can be misleading indicators of content quality (Wu et al., 2017; Yi et al., 2014). For instance, content with an eye-catching title but poor content quality may initially draw attention and obtain positive feedback, but only to later abuse the users’ trust and undermine their loyalty. This highlights a pressing need to balance the RS’s focus between short-term user engagement and long-term user satisfaction (Viljanen et al., 2016) . To this end, methods based on Reinforcement Learning (RL) have emerged (Chen et al., 2021b). By modeling sequential user behaviors using a Markov Decision Process (MDP), RL-based systems can dynamically adapt recommendations at every user interaction juncture (Zhao et al., 2011; Mahmood and Ricci, 2007). The strength of RL lies in its capability to optimize cumulative rewards over extended sequences in the future, allowing it to capture long-term objectives and trends, and ensure sustained user engagement. A recent finding (Cai et al., 2023a) shows that RL-based RS can even directly optimize user retention, a crucial metric often ignored but paramount in real-world business contexts (Zou et al., 2019; Chen et al., 2023) that is closely related to daily active users (DAU).
@@
  • Our innovative framework applies a novel multi-reward setting that balances immediate feedback with long-term retention by user-specific features, and then complements by a corresponding high-dimensional embedding module and a contrastive loss term.
@@ 3.2. Framework Overview
  • Embedding Module: The modified trajectory data then undergoes an embedding transformation to be represented as dense vectors. To be specific, the state and action sequences are processed using a Gated Recurrent Unit (GRU) (Cho et al., 2014), detailed further in Appendix B.1. In contrast, the reward sequence employs a high-dimensional encoder which is further explained in Section 3.4. Additionally, a positional session encoding is integrated.
  • Transformer Decision Block: Upon processing, the dense trajectory vectors serve as context, guiding the generation of action embeddings for the subsequent timestamp. This decision-making module is underpinned by the transformer model, detailed further in the Appendix B.2.
  • Action Decoding Block: Using the predicted action embeddings, our decoding unit strives to construct an action sequence that aligns closely with the reference of ground truth actions, detailed further in Appendix B.3.
@@ 3.3. Adaptive RTG Balancing
In settings that involve long short-term reward optimization, striking a balance between short-term and long-term performance metrics is notably challenging. This complexity arises from inherent conflicts between immediate and deferred objectives, as well as from the convoluted interdependencies among tasks (Ratner et al., 2018). To adeptly balance this equilibrium, we propose a novel strategy: the balancing of the RTG sequence, adaptively controlled by distinct user features. Since these user features hold a wealth of information about user preferences, they serve as a reliable guide for understanding user behaviors, bridging the gap between their immediate feedback and long-term retention.
@@
    where $z_{i}$ represents the $i$-th element of the output vector $\mathbf{z}$, and $K$ is the total number of elements in $\mathbf{z}$, representing weights for balancing immediate response and long-term retention.
@@ 3.4. Multi-reward Embedding
    (11) $\widehat{\textbf{E}}^{R}=concate(w_{1}E^{R}_{s,t},w_{2}E^{R}_{I,t})$
@@ 4.1. Dataset
We carried out our experiment using three datasets.
@@ 4.3. Baselines
  • RMTL (Liu et al., 2023b): This is a reinforcement learning-based multi-task recommendation model with adaptive loss weights.
@@
Furthermore, we’ve adapted the model architectures of MMoE and PLE to facilitate sequential recommendations, utilizing a weighted score derived from multiple tasks.
@@ 4.5. Overall Performance and Comparison
  • Superiority of DT4IER: The DT4IER model consistently outperforms the four baselines in both the realms of recommendation accuracy and user retention score across datasets. Particularly on the Kuairand-Pure dataset, DT4IER exhibits 0.07-0.09 improvement in recommendation accuracy compared to the top-performing baseline. By interpreting RL as an autoregressive setting and integrating a structure designed for multiple rewards, our model strikes a balance between immediate feedback and long-term retention, resulting in significant enhancements in recommendation performance while retaining user engagement.
@@ 4.8. Case Study
In this subsection, we highlight DT4IER’s effectiveness in improving recommendation performance through a case study on the Kuairand-Pure dataset. The selected user for this study has a history of four interactions, with the ground truth action sequence comprising details of five different videos. Figure 2 shows that when provided with the user’s state and the goal of maximizing retention, DT4Rec recommends four videos. This recommendation achieves a BLEU score of 0.625 and an NDCG of 0.710. In contrast, DT4IER outperforms by recommending five actions, all of which align with the ground truth action sequence. This leads to a higher prediction accuracy, with a BLEU score of 0.887 and an NDCG of 0.947. These results underscore our model’s enhanced efficiency, attributed to its consideration of short-term clicks and long-term retention.
@@ 5.1. Sequential Recommender Systems
Sequential Recommendation refers to a recommendation system paradigm that models patterns of user behavior and items over time to suggest relevant products or content (Li et al., 2022; Zhang et al., 2024; Gao et al., 2024; Liu et al., 2023c, a). Among the myriad of approaches available, Markov Chains (MCs) (Rendle et al., 2010) and Recurrent Neural Networks (RNNs) stand out for their prowess in sequence modeling (Donkers et al., 2017). The adaptability and capability of RNNs in integrating diverse information make them especially effective for crafting sequential recommendations. RU4Rec (Tan et al., 2016) was a pioneer in leveraging RNNs, tapping into their potential to process sequential data. Subsequently, GRU4Rec (Hidasi et al., 2016) unveiled a parallel RNN structure to process item features, thus boosting recommendation quality. Further, the generalization power of the transformer architecture has given rise to its prominence in Sequential Recommendation Systems, birthing models such as BERT4Rec (Sun et al., 2019) and SASRec (Kang and McAuley, 2018). Building upon these foundations, the decision transformer was conceptualized to tackle user retention challenges, aiming to directly predict actions using a reward-driven autoregressive framework (Zhao et al., 2023b). However, a gap remains in current research concerning the optimization of multi-reward settings.
@@ 5.2. Reinforcement Learning Based Recommender Systems
The validity of applying RL-based solutions (Sutton and Barto, 2018; Afsar et al., 2021; Wang et al., 2022; Zhang et al., 2022; Zhao et al., 2018a, 2021) for sequential recommendation comes from the assumption of Markov Decision Process (Shani et al., 2005). And the key advantage of RL solutions is the ability to improve the expected cumulative reward of future interactions with users, rather than optimizing the one-step recommendation. Specifically, for scenarios with small recommendation spaces, one can use tabular-based (Mahmood and Ricci, 2007; Moling et al., 2012) or value-based methods  (Taghipour et al., 2007; Zheng et al., 2018; Zhao et al., 2018b; Ie et al., 2019b) to directly evaluate the long-term value of the recommendation; For scenarios where action spaces are large, policy gradient methods (Sun and Zhang, 2018; Chen et al., 2019a, b; Zhao et al., 2020b) and actor-critic methods (Sutton et al., 1999; Peters and Schaal, 2008; Bhatnagar et al., 2007; Degris et al., 2012; Dulac-Arnold et al., 2015; Liu et al., 2018, 2020; Xin et al., 2020; Zhao et al., 2020a; Cai et al., 2023b; Liu et al., 2024) are adopted to guide the policy towards better recommendation quality. Knowing that the web service may want to optimize multiple metrics, several works have discussed the challenge of multi-objective optimization (Chen et al., 2021c; Cai et al., 2023b) where the user behaviors might have different distributional patterns. Among all user feedback signals, user retention has been considered one of the most challenging to optimize, while recent work has shown a possible RL-based approach (Cai et al., 2023a) for this uphill struggle. This work aims to simultaneously optimize immediate feedback and user retention. Similar to our work, to overcome the gap between experiments on real user environments and offline evaluations, user simulators are widely used to bypass this paradox (Ie et al., 2019c; Zhao et al., 2019, 2023a). Our approach can be thought of as a likelihood-based method while employing a sequence modeling objective rather than relying on variational techniques.
@@ 5.3. Multi-task Learning in Recommender Systems
Multi-task learning (MTL) is a machine learning technique that addresses multiple tasks simultaneously  (Zhang and Yang, 2021; Wang et al., 2023). It captures a shared representation of the input through a shared bottom layer and then processes each task using distinct networks with task-specific weights. The overall performance is further enhanced by the knowledge transfer between tasks. This approach has become particularly popular in recommender systems, thanks to its prowess in effectively sharing data across tasks and in recognizing a range of user behaviors (Ma et al., 2018a; Lu et al., 2018; Hadash et al., 2018; Pan et al., 2019; Pei et al., 2019). A significant portion of recent research targets the enhancement of these architectures to promote more effective knowledge sharing. Notably, some innovations focus on introducing constraints to task-specific parameters (Duong et al., 2015; Misra et al., 2016; Yang and Hospedales, 2016), while others aim to clearly demarcate shared from task-specific parameters (Ma et al., 2018b; Tang et al., 2020). The principal goal of these strategies is to amplify knowledge transfer through improved feature representation. Additionally, some researchers are exploring the potential of reinforcement learning to enhance the MTL model by adjusting loss function weights (Liu et al., 2023b). While many approaches prioritize item-wise modeling, our research delves into the challenges of balancing both short-term and long-term rewards in sequential recommendation settings. Furthermore, we have refined the methodologies of both the PLE and MMoE models for sequential recommendations.
@@ Appendix A HYPER-PARAMETER SELECTION
Appendix B Other Model details
B.1. State-action Encoder